# Learning Generalizable Features for Tibial Plateau Fracture Segmentation Using Masked Autoencoder and Limited Annotations

### v1

@@ Abstract
Accurate automated segmentation of tibial plateau fractures (TPF) from computed tomography (CT) requires large amounts of annotated data to train deep learning models, but obtaining such annotations presents unique challenges. The process demands expert knowledge to identify diverse fracture patterns, assess severity, and account for individual anatomical variations, making the annotation process highly time-consuming and expensive. Although semi-supervised learning methods can utilize unlabeled data, existing approaches often struggle with the complexity and variability of fracture morphologies, as well as limited generalizability across datasets. To tackle these issues, we propose an effective training strategy based on masked autoencoder (MAE) for the accurate TPF segmentation in CT. Our method leverages MAE pretraining to capture global skeletal structures and fine-grained fracture details from unlabeled data, followed by fine-tuning with a small set of labeled data. This strategy reduces the dependence on extensive annotations while enhancing the model’s ability to learn generalizable and transferable features. The proposed method is evaluated on an in-house dataset containing 180 CT scans with TPF. Experimental results demonstrate that our method consistently outperforms semi-supervised methods, achieving an average Dice similarity coefficient (DSC) of 95.81%, average symmetric surface distance (ASSD) of 1.91mm, and Hausdorff distance (95HD) of 9.42mm with only 20 annotated cases. Moreover, our method exhibits strong transferability when applying to another public pelvic CT dataset with hip fractures, highlighting its potential for broader applications in fracture segmentation tasks. The code will be publicly available.
@@ I Introduction
Semi-supervised learning, which utilizes a small set of labeled images in combination with a large pool of unlabeled images, has been successfully applied to various bone segmentation tasks [2, 3, 4, 5, 6, 7, 8, 9]. However, these semi-supervised learning methods still face limitations in TPF segmentation tasks. Firstly, existing semi-supervised learning methods for bone segmentation often rely on consistency regularization or self-training approaches [10, 11, 12, 13, 14, 15]. These methods typically use predictions generated by a teacher model on unlabeled data to supervise the training of a student model. However, as shown in Fig. 1, the diverse types and complex fragment morphologies of TPF pose huge challenges. Models trained on a small set of finely annotated data struggle to generate high-quality predictions for these highly variable fracture regions, ultimately affecting the overall training performance. Secondly, models trained using semi-supervised learning methods often exhibit satisfactory performance only on specific datasets, making it difficult to generalize or transfer their effectiveness to other datasets.

### v2

@@ Abstract
Accurate automated segmentation of tibial plateau fractures (TPF) from computed tomography (CT) requires large amounts of annotated data to train deep learning models, but obtaining such annotations presents unique challenges. The process demands expert knowledge to identify diverse fracture patterns, assess severity, and account for individual anatomical variations, making the annotation process highly time-consuming and expensive. Although semi-supervised learning methods can utilize unlabeled data, existing approaches often struggle with the complexity and variability of fracture morphologies, as well as limited generalizability across datasets. To tackle these issues, we propose an effective training strategy based on masked autoencoder (MAE) for the accurate TPF segmentation in CT. Our method leverages MAE pretraining to capture global skeletal structures and fine-grained fracture details from unlabeled data, followed by fine-tuning with a small set of labeled data. This strategy reduces the dependence on extensive annotations while enhancing the model’s ability to learn generalizable and transferable features. The proposed method is evaluated on an in-house dataset containing 180 CT scans with TPF. Experimental results demonstrate that our method consistently outperforms semi-supervised methods, achieving an average Dice similarity coefficient (DSC) of 95.81%, average symmetric surface distance (ASSD) of 1.91mm, and Hausdorff distance (95HD) of 9.42mm with only 20 annotated cases. Moreover, our method exhibits strong transferability when applying to another public pelvic CT dataset with hip fractures, highlighting its potential for broader applications in fracture segmentation tasks. Our code is publicly available at https://github.com/yuepeiyan/TPFSeg-MAE.
@@ I Introduction
Semi-supervised learning, which utilizes a small set of labeled images in combination with a large pool of unlabeled images, has been successfully applied to various bone segmentation tasks [2, 3, 4, 5, 6, 7]. However, these semi-supervised learning methods still face limitations in TPF segmentation tasks. Firstly, existing semi-supervised learning methods for bone segmentation often rely on consistency regularization or self-training approaches [8, 9, 10, 11, 12, 13, 14, 15]. These methods typically use predictions generated by a teacher model on unlabeled data to supervise the training of a student model. However, as shown in Fig. 1, the diverse types and complex fragment morphologies of TPF pose huge challenges. Models trained on a small set of finely annotated data struggle to generate high-quality predictions for these highly variable fracture regions, ultimately affecting the overall training performance. Secondly, models trained using semi-supervised learning methods often exhibit satisfactory performance only on specific datasets, making it difficult to generalize or transfer their effectiveness to other datasets.